OpenStudy (anonymous):

is 3.2-4d=2.3d+3 no sol. or identity

OpenStudy (anonymous):

show work plz

OpenStudy (anonymous):

nope a decimal

OpenStudy (anonymous):

-4d-2.3d=3-3.2 -6.3d=-0.2 6.3d=0.2 d=0.2/6.3

OpenStudy (anonymous):

thx